16. Find the solution to the following system of equations:
2x+3y=7
-x+4y=2

Answer by Cromlix(4381)

2x+3y=7 ...1
-x+4y=2 ...2
Multiply (2) by 2
2x + 3y = 7 ....1
-2x + 8y = 4 ....3
Add (1) + (3)
11y = 11
y = 1
Substitute y = 1 in (1)
2x + 3y = 7
2x + 3(1) = 7
2x + 3 = 7
2x = 7 - 3
2x = 4
x = 2
Your solution - {2,1}
Hope this helps
:-)